# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

### v3

@@ 4 Algorithmic Framework
Algorithm 2 proceeds in epochs, and each epoch $e$ contains multiple rounds of RL agent’s interaction with the environment. At each $t$, by playing action $a_{t}$ for insantaneous state $s_{t}$ through current epoch policy $\pi_{e}$, the agent collects next state configuration consisting of classical signals $s_{t+1},r(s_{t},a_{t})$ and a quantum sample $|\bm{\psi}\rangle^{(t)}$. Additionally, the agent keeps track of its visitations to the different states via the variables $\{\nu_{s,a},\mu_{s,a,s^{\prime}}\}$. Note that a new epoch is triggered at the “if” block of Line 15 in algorithm 2, which in turn is aligned to the widely used concept of doubling trick used in model-based RL algorithm design. Finally, the empirical visitation counts and probabilities up to epoch $e$, i.e., $\{N_{e+1}(s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ attributes are updated (line 15-21), and the policy is updated via re-solving from Eq. (15) - (19) and using Eq. (26).
5 Theoretical Results
In this section, we characterize the cumulative regret for $T$ rounds by running Q-UCRL in a unknown MDP hybrid Q-Environment. In the following, we first present a crucial result bounding the error accumulated till $e$ epochs between the true transition probabilities $\{P(\cdot|s,a)\}$ and the agent’s estimates $\{P_{e}(\cdot|s,a)\}$.
@@ 5 Theoretical Results
$\mathbb{P}[|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}]\geq 1-\delta$ (27)
@@
$\mathbb{P}[{||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{\log(S/\delta)}{n_{e+1}^{s_{i},a_{i}}}}]\geq 1-\delta,$ (28)
Thus, by substituting $n_{e}^{s_{i},a_{i}}$ with $\nu_{e}{(s,a)}/C$, Eq. (28) would produce:
${\mathbb{P}[||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{C\log(S/\delta)}{\nu_{e}(s,a)}]\geq 1-\delta}$ (29)
@@
Lemma 4.
@@
Here, it is critical to highlight that we obtain this high probability bound for model errors by careful utilization of Lemma 1 and quantum signals $\bm{\psi}^{t}$ via Eq. (42) - (49), while avoiding classical concentration bounds for probability norms that use martingale style analysis such as in the work (?). Finally, plugging the bound of term (a) as described by event $\mathcal{E}$, we obtain the following with probability $1-1/t_{e}^{6}$:
@@
Remark 2.
Note that for probability vectors ${P}_{e}(\cdot|s,a),{P}(\cdot|s,a)$ we have the bound: $\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1}\leq 2S$, in the worst case scenario. However, the probability that a bad event (i.e., complementary of Eq. (42)) happens is at most $\frac{1}{t_{e}^{6}}$.
Regret Decomposition: Regret of Q-UCRL defined in Eq. (10) can be decomposed as follows:
$\mathcal{R}_{[0,T-1]}$ $=T.\Gamma_{\pi^{*}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (54)
$=T.\Gamma_{\pi^{*}}^{P}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (55)
$=\sum_{e=1}^{E}T_{e}\big{(}\underbrace{\Gamma_{\pi^{*}}^{P}-\Gamma_{\pi^{*}_{e}}^{P}}_{\text{(a)}}\big{)}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (56)
@@
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (59)
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}$
$+\Big{[}\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\underbrace{\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}[r(s_{t},a_{t})]}_{\text{(}b)}\Big{]},$ (60)
@@
which essentially captures the gap in the average rewards by playing action $a$ in state $s$ accordingly to some policy $\pi$ in one step of the optimistic transition model $P_{e}$ w.r.t. the actual probabilities $P$. Also, we note here that the definition of $Q_{\pi}^{{P}_{e}}$ function is associated to the value function definition in eq (5).
This leads us to the following key results that connect the regret expression in Eq. (61) and the Bellman errors.
@@
[ Lemma 5.2, (?)] The difference in long-term expected reward by playing optimistic policy of epoch $e$, i.e., $\pi_{e}$ on the optimistic transition model estimates, i.e., $\Gamma_{\pi_{e}}^{{P}_{e}}$, and the expected reward collected by playing $\pi_{e}$ on the actual model $P$, i.e., $\Gamma_{\pi_{e}}^{{P}}$ is the long-term average of the Bellman errors. Mathematically, we have:
$\Gamma_{\pi_{e}}^{{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}=\sum_{s,a}\rho_{\pi_{e}}^{P}(s,a)B_{\pi_{e}}^{{P}_{e}}(s,a),$ (63)
where $B_{\pi}^{{P}_{e}}(s,a)$ is defined in Eq. (62).
With Eq. (62), it necessary to mathematically bound the Bellman errors. To this end, we leverage the following result that characterizes the Bellman errors in terms of $\tilde{h}(\cdot)$, a quantity that measures the bias contained in the optimistic MDP model.
Lemma 6.
[Lemma 5.3, (?)] The Bellman errors for any arbitrary state-action pair $(s,a)$ corresponding to epoch $e$ is bounded as follows:
where $\tilde{h}(\cdot)$ (defined in (?)) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval.
Consequently, by plugging in the bound of the error arising from estimation of transition model in Lemma 4 into eq 64, we obtain the following result with probability $1-1/t_{e}^{6}$:
$B^{\pi_{e},{P}_{e}}(s,a)$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}\|\tilde{h}(\cdot)\|_{\infty},$ (65)
Consequently, it is necessary to bound the model bias error term in Eq. (65). We recall that the optimistic MDP $P_{e}$ corresponding to epoch $e$ maximizes the average reward in the confidence set as described in the formulation of eq (15). In other words, $\Gamma_{\pi_{e}}^{P_{e}}\geq\Gamma_{\pi_{e}}^{P^{\prime}}$ for every $P^{\prime}$ in the confidence interval. Specifically, the model bias for optimistic MDP can be explicitly defined in terms of Bellman equation (?), as follows:
where the following simplified notations are used: $r_{\pi_{e}}(s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)r(s,a)$, $P_{e,\pi_{e}}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)P_{e}(s^{\prime}|s,a)$. Next, we incorporate the following result that correlates model bias $\tilde{h}(\cdot)$ to the mixing time $T_{\texttt{mix}}$ of MDP $\mathcal{M}$.
Lemma 7.
[Lemma D.5, (?)] For an MDP with the transition model $P_{e}$ that generates rewards $r(s,a)$ on playing policy $\pi_{e}$, the difference of biases for states $s$ and $s^{\prime}$ are bounded as:
@@
Plugging in the bound of bias term $\tilde{h}(\cdot)$ into Eq. (65), we get the final bound for Bellman error with probability $1-1/T^{6}$ as:
$B^{\pi_{e},{P}_{e}}(s,a)$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}.$ (68)
Regret Bound for Q-UCRL We now present the final bound for regret incurred by Q-UCRL over a $T$ length training horizon.
Theorem 1.

### v4

@@ Introduction
Algorithm 2 proceeds in epochs, and each epoch $e$ contains multiple rounds of RL agent’s interaction with the environment. At each $t$, by playing action $a_{t}$ for insantaneous state $s_{t}$ through current epoch policy $\pi_{e}$, the agent collects next state configuration consisting of classical signals $s_{t+1},r(s_{t},a_{t})$ and a quantum sample $|\bf{\psi}\rangle^{(t)}$. Additionally, the agent keeps track of its visitations to the different states via the variables $\{\nu_{s,a},\mu_{s,a,s^{\prime}}\}$. Note that a new epoch is triggered at the “if” block of Line 15 in algorithm 2, which in turn is aligned to the widely used concept of doubling trick used in model-based RL algorithm design. Finally, the empirical visitation counts and probabilities up to epoch $e$, i.e., $\{N_{e+1}(s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ attributes are updated (line 15-21), and the policy is updated via re-solving from Eq. (15) - (19) and using Eq. (26).
Theoretical Results
In this section, we characterize the cumulative regret for $T$ rounds by running Q-UCRL in a unknown MDP hybrid Q-Environment. In the following, we first present a crucial result bounding the error accumulated till $e$ epochs between the true transition probabilities $\{P(\cdot|s,a)\}$ and the agent’s estimates $\{\hat{P}_{e}(\cdot|s,a)\}$.
@@
$\mathbb{P}\bigg{[}|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}\bigg{]}\geq 1-\delta$ (27)
@@
$\mathbb{P}\bigg{[}{||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{\log(S/\delta)}{n_{e+1}^{s_{i},a_{i}}}}\bigg{]}\geq 1-\delta,$ (28)
By switching $n_{e}^{s_{i},a_{i}}$ with $\nu_{e}{(s,a)}/C$, Eq. (28) gives:
${\mathbb{P}\bigg{[}||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{C\log(S/\delta)}{\nu_{e}(s,a)}\bigg{]}\geq 1-\delta}$ (29)
@@
Note that the choice of $C$ in Eq. (27) is to ensure that the samples $\nu_{e}(s,a)$ are sufficient for performing Algorithm 1.
@@
$\mathbb{P}\bigg{[}|\hat{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{eC\log(e{S}/\delta)}{N_{e+1}(s,a)}\bigg{]}\geq 1-\delta$ (30)
Proof.
Please refer to Appendix C. ∎
Lemma 2 and Lemma 3 provide the acceleration for mean estimation, which further improve global convergence rate.
Lemma 4.
$\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1},$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}},$ (31)
Proof.
Please refer to Appendix D. ∎
Remark 1.
Note that for ${P}_{e}(\cdot|s,a),{P}(\cdot|s,a)$ we have the bound: $\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1}\leq 2S$, in the worst case scenario. However, the probability that a bad event (i.e., complementary of Eq. (31)) happens is at most $\frac{1}{t_{e}^{6}}$.
Lemma 5 (Regret Decomposition).
Regret of Q-UCRL defined in Eq. (10) can be decomposed as follows:
$\mathcal{R}_{[0,T-1]}\leq\sum_{e=1}^{E}T_{e}\Big{(}\Gamma_{\pi_{e}}^{\tilde{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}\Big{)}.$ (32)
Proof.
Please refer to Appendix E. ∎
which essentially captures the gap in the average rewards by playing action $a$ in state $s$ accordingly to some policy $\pi$ in one step of the optimistic transition model $P_{e}$ w.r.t. the actual probabilities $P$. This leads us to the next result that ties the regret expression in Eq. (32) and the Bellman errors.
Lemma 6 (Lemma 5.2, (agarwal2021concave)).
The difference in long-term expected reward by playing optimistic policy of epoch $e$, i.e., $\pi_{e}$ on the optimistic transition model estimates, i.e., $\Gamma_{\pi_{e}}^{{P}_{e}}$, and the expected reward collected by playing $\pi_{e}$ on the actual model $P$, i.e., $\Gamma_{\pi_{e}}^{{P}}$ is the long-term average of the Bellman errors. Mathematically, we have:
$\Gamma_{\pi_{e}}^{\tilde{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}=\sum_{s,a}\rho_{\pi_{e}}^{P}(s,a)B_{\pi_{e}}^{{P}_{e}}(s,a),$ (34)
where $B_{\pi}^{{P}_{e}}(s,a)$ is defined in Eq. (33).
With Eq. (34), it necessary to mathematically bound the Bellman errors. To this end, we leverage the following result from (agarwal2021concave) which is stated as Lemma 10 from Appendix F characterizing the Bellman errors in terms of $\tilde{h}(\cdot)$, a quantity that measures the bias contained in the optimistic MDP model.
where $\tilde{h}(\cdot)$ (defined in (puterman2014markov)) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval. Using the bound of the error arising from estimation of transition model in Lemma 4 into Eq. (35), we obtain the following result with probability $1-1/t_{e}^{6}$:
$B^{\pi_{e},{P}_{e}}(s,a)$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCe\log{(eS)}}{\max\{1,N_{e}(s,a)\}}\|\tilde{h}(\cdot)\|_{\infty},$ (36)
Consequently, it is necessary to bound the model bias error term in Eq. (36). We recall that the optimistic MDP $P_{e}$ corresponding to epoch $e$ maximizes the average reward in the confidence set as described in the formulation of Eq. (15). In other words, $\Gamma_{\pi_{e}}^{P_{e}}\geq\Gamma_{\pi_{e}}^{P^{\prime}}$ for every $P^{\prime}$ in the confidence interval. Specifically, the model bias for optimistic MDP can be explicitly defined in terms of Bellman equation (puterman2014markov), as follows:
where the following simplified notations are used: $r_{\pi_{e}}(s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)r(s,a)$, $P_{e,\pi_{e}}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)P_{e}(s^{\prime}|s,a)$. Next, we incorporate the following result from (agarwal2021concave) stated as Lemma 11 in Appendix F correlating model bias $\tilde{h}(\cdot)$ to the mixing time $T_{\texttt{mix}}$ of MDP $\mathcal{M}$.
Using bound of bias term $\tilde{h}(\cdot)$ into Eq. (36), we get the final bound for Bellman error with probability $1-1/T^{6}$ as:
$\small B^{\pi_{e},{P}_{e}}(s,a)\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCe\log{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}.$ (39)
Note that Q-UCRL obtains a quadratic improvement for the Bellman error over the classical UC-UCRL (agarwal2021concave). We next present the final regret bound of Q-UCRL.
Theorem 1.
$\small\mathcal{R}_{[0,T-1]}=\mathcal{O}\Bigg{(}\textstyle{{\!S^{5}A^{4}T_{\texttt{mix}}{\log^{3}\bigg{(}\!\frac{T}{SA}\!\bigg{)}\log^{\frac{1}{2}}(T\sqrt{S})\log(S^{2}AT)}}}\!\!\Bigg{)}.$ (40)
Proof.
Please refer to Appendix G. ∎
Remark 2.
We reiterate that our characterization of regret incurred by Q-UCRL in Theorem 1 is a martingale-free analysis and thus deviates from its classical counterparts (fruit2018efficient; auer2008near; agarwal2021concave).
Conclusions
A promising future direction is parameterized quantum RL, which has recently shown speedups for discounted settings (xu2025accelerating); achieving similar gains for average-reward problems, where the state-of-the-art in non-quantum RL is given by (ganesh2025sharper), remains an open challenge.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Reproducibility Checklist
This paper
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes)
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes)
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper (yes)
Does this paper make theoretical contributions? (yes)
If yes, please complete the list below.
  • All assumptions and restrictions are stated clearly and formally. (yes)
  • All novel claims are stated formally (e.g., in theorem statements). (yes)
  • Proofs of all novel claims are included. (yes) Proof sketches or intuitions are given for complex and/or novel results. (yes)
  • Appropriate citations to theoretical tools used are given. (yes)
  • All theoretical claims are demonstrated empirically to hold. (no)
  • All experimental code used to eliminate or disprove claims is included. (NA)
Does this paper rely on one or more datasets? (no)
If yes, please complete the list below.
  • A motivation is given for why the experiments are conducted on the selected datasets (yes)
  • All novel datasets introduced in this paper are included in a data appendix. (yes/partial/no/NA)
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no/NA)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes/no/NA)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes/partial/no/NA)
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (yes/partial/no/NA)
Does this paper include computational experiments? (no)
If yes, please complete the list below.
  • Any code required for pre-processing data is included in the appendix. (yes/partial/no).
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (yes/partial/no)
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes/partial/no)
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no)
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (yes/partial/no/NA)
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes/partial/no)
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes/partial/no)
  • This paper states the number of algorithm runs used to compute each reported result. (yes/no)
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes/no)
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes/partial/no)
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA)
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA)
Appendix A Appendix A: Convexity of Optimization Problem
In order to solve the optimistic optimization problem in Eq. (15) - (19) for epoch $e$, we adopt the approach proposed in [agarwal2021concave, rosenberg2019online]. The key is that the constraint in (18) seems non-convex. However, that can be made convex with an inequality as
Appendix B Appendix B: A brief overview of Quantum Amplitude Amplification
  1. 1.
    Quantum Search: In the quantum algorithm for unstructured search, known as Grover’s algorithm [grover1996fast], the technique of amplitude amplification is employed to enhance the amplitude of the target state and decrease the amplitude of the non-target states. As a result, this technique provides a quadratic improvement over classical search algorithms.
  2. 2.
    Quantum Optimization: Quantum amplification can be used in quantum optimization algorithms, such as Quantum Approximate Optimization Algorithm (QAOA) [farhi2014quantum], to amplify the amplitude of the optimal solution and reduce the amplitude of sub-optimal solutions. This can lead to an exponential speedup in solving combinatorial optimization problems.
  3. 3.
    Quantum Simulation: Quantum amplification has also been used in quantum simulation algorithms, such as Quantum Phase Estimation (QPE) [kitaev1995quantum], to amplify the amplitude of the eigenstate of interest and reduce the amplitude of other states. This can lead to an efficient simulation of quantum systems, which is an intractable problem in the classical domain.
  4. 4.
    Quantum Machine Learning: The utilization of quantum amplification is also evident in quantum machine learning algorithms, including Quantum Support Vector Machines (QSVM) [lloyd2013quantum]. The primary objective is to amplify the amplitude of support vectors while decreasing the amplitude of non-support vectors, which may result in an exponential improvement in resolving particular classification problems.
In the domain of quantum Monte Carlo methods, implementing quantum amplitude amplification can enhance the algorithm’s efficiency by decreasing the amount of samples needed to compute the integral or solve the optimization problem. By amplifying the amplitude of the desired state, a substantial reduction in the number of required samples can be achieved while still maintaining a certain level of accuracy, leading to significant improvements in efficiency when compared to classical methods. This technique has been particularly useful in [hamoudi2021quantum] for achieving efficient convergence rates in quantum Monte Carlo simulations.
Appendix C Appendix C: Proof of Lemma 3
Lemma 7 (Lemma 3 in the main paper).
@@ Appendix C Appendix C: Proof of Lemma 3
Appendix D Appendix D: Proof of Lemma 4
Lemma 8 (Lemma 4 in the main paper).
@@ Appendix D Appendix D: Proof of Lemma 4
Here, it is critical to highlight that we obtain this high probability bound for model errors by careful utilization of Lemma 1 and quantum signals $|\bf{\psi}\rangle^{t}$ via Eq. (55) - (62), while avoiding classical concentration bounds for probability norms that use martingale style analysis such as in the work [weissman2003inequalities]. Finally, plugging the bound of term (a) as described by event $\mathcal{E}$, we obtain the following with probability $1-1/t_{e}^{6}$:
@@
Appendix E Appendix E: Proof of Lemma 5
Lemma 9 (Regret Decomposition, Same as Lemma 5 in the main paper).
Regret of Q-UCRL defined in Eq. (10) can be decomposed as follows:
$\mathcal{R}_{[0,T-1]}\leq\sum_{e=1}^{E}T_{e}\Big{(}\Gamma_{\pi_{e}}^{\tilde{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}\Big{)}.$ (67)
Proof.
$\mathcal{R}_{[0,T-1]}=T\times\Gamma_{\pi^{*}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (68)
$=T\times\Gamma_{\pi^{*}}^{P}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$
$=\sum_{e=1}^{E}T_{e}\big{(}\underbrace{\Gamma_{\pi^{*}}^{P}-\Gamma_{\pi^{*}_{e}}^{P}}_{\text{(a)}}\big{)}+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (69)
@@ Appendix E Appendix E: Proof of Lemma 5
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}+\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (72)
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}+\Big{[}\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\underbrace{\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}[r(s_{t},a_{t})]}_{\text{(}b)}\Big{]},$ (73)
@@
Appendix F Appendix F: Some Auxiliary Lemmas used in this Paper
Lemma 10 (Lemma 5.3, [agarwal2021concave]).
The Bellman errors for any arbitrary state-action pair $(s,a)$ corresponding to epoch $e$ is bounded as follows:
where $\tilde{h}(\cdot)$ (defined in [puterman2014markov]) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval.
@@ Appendix F Appendix F: Some Auxiliary Lemmas used in this Paper
[Lemma D.5, [agarwal2021concave]] For an MDP with the transition model $P_{e}$ that generates rewards $r(s,a)$ on playing policy $\pi_{e}$, the difference of biases for states $s$ and $s^{\prime}$ are bounded as:
@@
Appendix G Appendix G: Proof of Theorem 1
Theorem 2 (Theorem 1 in main paper).